Consider an isolated system that contains two pieces of copper separated by an (internal) insulating wall. Initially, the first piece is at 500K and the second is at 300K. Calculate the entropy change in the system when the insulating wall is removed; assume that each piece has half a mole of copper in it. Given: Specific heat capacity of Cu is

CP = 22.6JK−1mol−1

. [Hint: S is an extensive state function].

(b) What is the entropy change of the universe ? Comment whether the process is reversible/irreversible ?

(c) Repeat your calculations for a molar heat capacity which is a function of T: CP = 22.6 + 6.28e − 03 T Jmol−1 K−1

How does the equilibrium T change for such a CP ? Can you justify it ?


1
Expert's answer
2020-12-24T12:17:49-0500

Answer

a) change in entropy

"\Delta S=\frac{nC_p dT}{T}=\frac{1\times22.6\times200}{500}"

="12.04" J/k


b) The second law of thermodynamics states that in a reversible process, the entropy of the universe is constant, whereas in an irreversible process, such as the transfer of heat from a hot object to a cold object, the entropy of the universe increases.

c) molar specific heat is given by

"C_p=\frac{dQ}{ndT}=22.6 + 6.28e^ {− 03 T }"

From above equation we can clearly say

When we increase T or temperature Cp

Start to decrease.